THE MAGELLANIC CLOUDS NEWSLETTER

An electronic publication dedicated to the Magellanic Clouds, and astrophysical phenomena therein

No. 114 — 1 December 2011

http://www.astro.keele.ac.uk/MCnews

Editor: Jacco van Loon

Figure 1: Model producing an object similar to SN 1987A (Zaninetti 2011; cf. MC News #113).

Editorial

Dear Colleagues,

It is my pleasure to present you the 114^{th} issue of the Magellanic Clouds Newsletter. This edition is remarkably voluminous, with no less than five refereed papers on Cepheids alone, many papers on massive stars including Be stars and X-ray binaries and several directly resulting from the VLT-FLAMES and/or Tarantula surveys. Likewise there are a couple of papers resulting from the VISTA Magellanic Clouds survey, the overview of the SAGE-SMC Spitzer survey and some interstellar medium results arising from Herschel and HST-COS surveys as well as improved radio continuum data. There are some interesting papers on the reverse shock in SN 1987A and a likely supernova progenitor (Sanduleak’s star), too.

Looking for a job? Check out the offers of postdoc positions in interstellar medium studies at STScI, or extra-galactic astrophysics at Toledo, both in the USA.

If you are looking for meetings, then check out the announcements of one on the mass loss return from stars to galaxies at STScI (USA), and another one on circumstellar dynamics at the border of Brazil with Argentina.

The next issue is planned to be distributed on the 1^{st} of February 2012. With the best wishes for the holiday season and a Happy New Year,

Editorially Yours, Jacco van Loon

Refereed Journal Papers

Calibrating the Cepheid Period–Luminosity relation from the infrared

surface brightness technique. II. The effect of metallicity, and the distance to the LMC

Jesper Storm^{1}, Wolfgang Gieren^{2}, Pascal Fouqué^{3}, Thomas G. Barnes^{4}, Igor Soszyński^{5}, Grzegorz Pietrzyński^{2,5}, Nicolas Nardetto^{6} and Didier Queloz^{7}

1Leibniz Institute for Astrophysics Potsdam (AIP), Germany 2Universidad de Concepción, Departamento Astronomía, Chile 3IRAP, Univ. de Toulouse, CNRS, France 4Univ. of Texas, McDonald Observatory, USA 5Warsaw University Observatory, Poland 6Laboratoire Fizeau, UNS/OCA/CNRS UMR6525, Nice, France 7Observatoire Astronomique de l’Université de Genève, Switzerland

The extragalactic distance scale builds directly on the Cepheid Period–Luminosity (PL) relation as delineated by the sample of Cepheids in the Large Magellanic Cloud (LMC). However, the LMC is a dwarf irregular galaxy, quite different from the massive spiral galaxies used for calibrating the extragalactic distance scale. Recent investigations suggest that not only the zero-point but also the slope of the Milky Way PL relation differ significantly from that of the LMC, casting doubts on the universality of the Cepheid PL relation. We want to make a differential comparison of the PL relations in the two galaxies by delineating the PL relations using the same method, the infrared surface brightness method (IRSB), and the same precepts. The IRSB method is a Baade–Wesselink type method to determine individual distances to Cepheids. We apply a newly revised calibration of the method as described in an accompanying paper (Paper I) to 36 LMC and five SMC Cepheids and delineate new PL relations in the V, I, J, & K bands as well as in the Wesenheit indices in the optical and near-IR. We present 509 new and accurate radial velocity measurements for a sample of 22 LMC Cepheids, enlarging our earlier sample of 14 stars to include 36 LMC Cepheids. The new calibration of the IRSB method is directly tied to the recent HST parallax measurements to ten Milky Way Cepheids, and we find a LMC barycenter distance modulus of 18.45 ± 0.04 (random error only) from the 36 individual LMC Cepheid distances. In the J, K bands we find identical slopes for the LMC and Milky Way PL relations and only a weak metallicity effect on the zero points (consistent with a zero effect), metal poor stars being fainter. In the optical we find the Milky Way slopes are slightly shallower than the LMC slopes (but again consistent with no difference in the slopes) and small effects on the zero points. However, the important Wesenheit index in V, (V − I) shows a metallicity effect on the slope and on the zero point which is likely to be significant. We find a significant metallicity effect on the W_{V I} index γ(W_{V I}) = −0.23 ± 0.10 mag dex^{−1} as well as an effect on the slope. The K-band PL relation on the other hand is found to be an excellent extragalactic standard candle being metallicity insensitive in both slope and zero-point and at the same time being reddening insensitive and showing the least internal dispersion.

Accepted for publication in Astronomy and Astrophysics Available from arXiv:1109.2016

The Baade–Wesselink p-factor applicable to LMC Cepheids

N. Nardetto^{1}, A. Fokin^{2}, P. Fouqué^{3}, J. Storm^{4}, W. Gieren^{5}, G. Pietrzyński^{5,6}, D. Mourard^{1} and P. Kervella^{7}

1Laboratoire Fizeau, UNS/OCA/CNRS UMR6525, Parc Valrose, 06108 Nice Cedex 2, France 2Institute of Astronomy of the Russian Academy of Sciences, 48 Pjatnitskaya Str., Moscow 109017 Russia 3Observatoire Midi-Pyrénées, Laboratoire d’Astrophysique, UMR 5572, Université Paul Sabatier – Toulouse 3, 14 avenue Edouard Belin, 31400 Toulouse, France 4Leibniz-Institut für Astrophysik Potsdam (AIP), An der Sternwarte 16, D-14482 Potsdam, Germany 5Departamento de Astronomía, Universidad de Concepción, Casilla 160-C, Concepción, Chile 6Warsaw University Observatory, Al. Ujazdowskie 4, 00-478, Warsaw, Poland 7LESIA, Observatoire de Paris, CNRSUMR8109, UPMC, Université Paris Diderot, 5 place Jules Janssen, 92195 Meudon, France

Context: Recent observations of LMC Cepheids bring new constraints on the slope of the period-projection factor relation (hereafter P p relation) that is currently used in the Baade–Wesselink (hereafter BW) method of distance determination. The discrepancy between observations and theoretical analysis is particularly significant for short period Cepheids. Aims: We investigate three physical effects that might possibly explain this discrepancy: (1) the spectroscopic S/N that is systematically lower for LMC Cepheids (around 10) compared to Galactic ones (up to 300), (2) the impact of the metallicity on the dynamical structure of LMC Cepheids, and (3) the combination of infrared photometry/interferometry with optical spectroscopy. Methods: To study the S/N we use a very simple toy model of Cepheids. The impact of metallicity on the projection factor is based on the hydrodynamical model of δ Cep already described in previous studies. This model is also used to derive the position of the optical versus infrared photospheric layers. Results: We find no significant effect of S/N, metallicity, and optical-versus-infrared observations on the P p relation. Conclusions: The P p relation of Cepheids in the LMC does not differ from the Galactic relation. This allows its universal application to determine distances to extragalactic Cepheids via BW analysis.

Accepted for publication in A&A Available from arXiv:1109.6763

The VLT-FLAMES survey of massive stars: Nitrogen abundances for Be-type stars in the Magellanic Clouds

P.R. Dunstall^{1}, I. Brott^{2}, P.L. Dufton^{1}, D.J. Lennon^{4}, C.J. Evans^{3}, S.J. Smartt^{1} and I. Hunter^{1}

1Department of Physics & Astronomy, The Queen’s University of Belfast, BT7 1NN, Northern Ireland, UK 2University of Vienna, Department of Astronomy, Türkenschanzstr. 17, A-1180, Vienna, Austria 3UK Astronomy Technology Centre, Royal Observatory Edinburgh, Blackford Hill, Edinburgh, EH9 3HJ, UK 4ESA, Space Telescope Science Institute, 3700 San Martin Drive, Baltimore, MD 21218, USA

We compare the predictions of evolutionary models for early-type stars with atmospheric parameters, projected rotational velocities and nitrogen abundances estimated for a sample of Be-type stars. Our targets are located in 4 fields centred on the Large Magellanic Cloud cluster: NGC 2004 and the N 11 region as well as the Small Magellanic Cloud clusters: NGC 330 and NGC 346. Atmospheric parameters and photospheric abundances have been determined using the non-LTE atmosphere code tlusty. Effective temperature estimates were deduced using three different methodologies depending on the spectral features observed; in general they were found to yield consistent estimates. Gravities were deduced from Balmer line profiles and microturbulences from the Si iii spectrum. Additionally the contributions of continuum emission from circumstellar discs were estimated. Given its importance in constraining stellar evolutionary models, nitrogen abundances (or upper limits) were deduced for all the stars analysed. Our nitrogen abundances are inconsistent with those predicted for targets spending most of their main sequence life rotating near to the critical velocity. This is consistent with the results we obtain from modelling the inferred rotational velocity distribution of our sample and of other investigators. We consider a number of possibilities to explain the nitrogen abundances and rotational velocities of our Be-type sample.

Accepted for publication in Astronomy & Astrophysics Available from arXiv:1109.6661

Non-standard grain properties, dark gas reservoir, and extended submillimeter excess, probed by Herschel in the Large Magellanic Cloud

Frédéric Galliano^{1}, Sacha Hony^{1}, Jean-Philippe Bernard^{2}, Caroline Bot^{3}, Suzanne C. Madden^{1}, Julia Roman-Duval^{4}, Maud Galametz^{5}, Aigen Li^{6}, Margaret Meixner^{4}, Charles W. Engelbracht^{7}, Vianney Lebouteiller^{1},

Karl Misselt^{7}, Edward Montiel^{7}, Pasquale Panuzzo^{1}, William T. Reach^{8} and Ramin Skibba^{7}

1AIM, CEA/Saclay, L’Orme des Merisiers, 91191 Gif-sur-Yvette, France 2Centre d’Étude Spatiale des Rayonnements, CNRS, 9 Av. du Colonel Roche, BP 4346, 31028 Toulouse, France 3Observatoire Astronomique de Strasbourg, 11 rue de l’université, 67000 Strasbourg, France 4Space Telescope Science Institute, 3700 San Martin Drive, Baltimore, MD 21218, USA 5Institute of Astronomy, University of Cambridge, Madingley Road, Cambridge CB3 0HA, UK 6314 Physics Building, Department of Physics and Astronomy, University of Missouri, Columbia, MO 65211, USA 7Steward Observatory, University of Arizona, 933 North Cherry Ave., Tucson, AZ 85721, USA 8Spitzer Science Center, California Institute of Technology, MS 220-6, Pasadena, CA 91125, USA

Context: Herschel provides crucial constraints on the IR SEDs of galaxies, allowing unprecedented accuracy on the dust mass estimates. However, these estimates rely on non-linear models and poorly-known optical properties. Aims: In this paper, we perform detailed modelling of the Spitzer and Herschel observations of the LMC, in order to: (i) systematically study the uncertainties and biases affecting dust mass estimates; and to (ii) explore the peculiar ISM properties of the LMC. Methods: o achieve these goals, we have modelled the spatially resolved SEDs with two alternate grain compositions, to study the impact of different submillimetre opacities on the dust mass. We have rigorously propagated the observational errors (noise and calibration) through the entire fitting process, in order to derive consistent parameter uncertainties. Results: First, we show that using the integrated SED leads to underestimating the dust mass by ≃ 50% compared to the value obtained with sufficient spatial resolution, for the region we studied. This might be the case, in general, for unresolved galaxies. Second, we show that Milky Way type grains produce higher gas-to-dust mass ratios than what seems possible according to the element abundances in the LMC. A spatial analysis shows that this dilemma is the result of an exceptional property: the grains of the LMC have on average a larger intrinsic submm opacity (emissivity index β ≃ 1.7 and opacity κ_{abs}(160μm)= 1.6 m^{2} kg^{−1}) than those of the Galaxy. By studying the spatial distribution of the gas-to-dust mass ratio, we are able to constrain the fraction of unseen gas mass between ≃ 10, and ≃ 100% and show that it is not sufficient to explain the gas-to-dust mass ratio obtained with Milky Way type grains. Finally, we confirm the detection of a 500 μm extended emission excess with an average relative amplitude of ≃ 15%, varying up to 40%. This excess anticorrelates well with the dust mass surface density. Although we do not know the origin of this excess, we show that it is unlikely the result of very cold dust, or CMB fluctuations. Accepted for publication in Astronomy & Astrophysics Available from arXiv:1110.1260

The temperature structure of Be star disks in the Small Magellanic Cloud

A. Ahmed^{1,2} and T.A.A. Sigut^{1}

1Department of Physics and Astronomy, The University of Western Ontario, London, Ontario, CANADA N6A 3K7 2Department of Astronomy, Cairo University, 12613 Giza, Egypt

The temperature structure of Be star circumstellar disks at the sub-solar metallicity appropriate to the Small Magellanic Cloud (SMC) is investigated. It is found that for central stars of the same spectral type, Be star disks in the SMC are systematically hotter by several thousand degrees compared to Milky Way (MW) disks with the same density structure. For early spectral types (B0e – B3e), this results in systematically smaller Hα equivalent widths for Be stars in the SMC. The implication of this result on Be star frequency comparisons between MW and SMC clusters is shown to be a 5 – 10% lowering of the detection efficiency of Be stars in SMC clusters. These calculations are also compared to the known Hα equivalent width distributions in the MW and SMC. For the MW, reasonable agreement is found; however, for the SMC, the match is not as good and systematically larger Be star disks may be required. Accepted for publication in ApJ Available from arXiv:1110.0547

Near-IR polarimetry around 30 Doradus: I. Separation of the Galactic sources

Jaeyeong Kim^{1}, Soojong Pak^{1}, Minho Choi^{2}, Wonseok Kang^{1}, Ryo Kandori^{3}, Motohide Tamura^{3}, Tetsuya Nagata^{4}, Jungmi Kwon^{3}, Daisuke Kato^{5} and Daniel T. Jaffe^{1,6}

1School of Space Research, Kyung Hee University, 1 Seocheon-dong, Giheung-gu, Yongin, Gyeonggi-do 446-701, South Korea 2Korea Astronomy and Space Science Institute, 776 Daedeokdaero, Yuseong, Daejeon 305-348, South Korea 3National Astronomical Observatory of Japan, Mitaka, Tokyo 181-8588, Japan 4Department of Astronomy, Kyoto University, Kyoto 606-8502, Japan 5Institute of Space and Astronomical Science, Japan Aerospace Exploration Agency, 3-1-1 Yoshinodai, Chuo-ku, Sagamihara, Kanagawa, 252-5210, Japan 6Department of Astronomy, University of Texas at Austin, 1 University Station, Austin, TX 78712-0259, USA

A 20′ × 20′ region around 30 Doradus in the Large Magellanic Cloud (LMC) is observed and analyzed in the nearinfrared. We obtain polarimetry data in the J, H, and K_{s} bands using the SIRIUS polarimeter SIRPOL at the Infrared Survey Facility 1.4 m telescope. We measure the Stokes parameters of 2562 point-like sources to derive the degree of polarization and the polarization position angles. We discuss the statistics of the groups classified by color–magnitude diagram and proper motions of the sources, in order to separate the Galactic foreground sources from those present in the LMC. We notice that groups classified by the proper motion data show a tendency towards different polarimetric properties.

Published in Journal of the Korean Astronomical Society, vol. 44, no. 4, p135 Available from arXiv:1110.4463

The VMC survey III. Mass-loss rates and luminosities of LMC AGB stars

M. Gullieuszik^{1}, M.A.T. Groenewegen^{1}, M.-R.L. Cioni^{2,3}, R. de Grijs^{4,5}, J.Th. van Loon^{6}, L. Girardi^{7}, V.D. Ivanov^{8}, J.M. Oliveira^{6}, J.P. Emerson^{9} and R. Guandalini^{2}

1Royal Observatory of Belgium, Belgium 2University of Hertforshire, United Kingdom 3University Observatory Munich, Germany 4Kavli Institute for Astronomy and Astrophysics, China 5Department of Astronomy and Space Science, Kyung Hee University, South Korea 6Keele University, United Kingdom 7INAF, Osservatorio Astronomico di Padova, Italy 8European Southern Observatory, Santiago, Chile 9Astronomy Unit, School of Physics & Astronomy, University of London, United Kingdom

Dust radiative transfer models are presented for all 374 AGB stars candidates in one of the fields observed by the new VISTA survey of the Magellanic Clouds (VMC). Mass-loss rates, luminosities and a classification of C- and O-rich stars are derived by fitting the models to the spectral energy distribution obtained by combining VMC data with existing optical, near-, and mid-infrared photometry. This exploratory study shows that our method provides reliable mass-loss rates, luminosities and chemical classifications for all AGB stars. These results offer already important constraints to AGB evolutionary models. Most of our conclusions, especially for the rarer dust-enshrouded extreme AGB stars, are however strongly limited by the relatively small area covered by our study. Forthcoming VMC observations will easily remove this limitation. [abridged]

Accepted for publication in Astronomy and Astrophysics Available from arXiv:1110.4497

Discovery of a Be/X-ray pulsar binary and associated supernova remnant in the Wing of the SMC

V. Hénault-Brunet^{1}, L.M. Oskinova^{2}, M.A. Guerrero^{3}, W. Sun^{4}, Y.-H. Chu^{5}, C.J. Evans^{6,1}, J.S. Gallagher III^{7}, R.A. Gruendl^{5} and J. Reyes-Iturbide^{8}

1Scottish Universities Physics Alliance (SUPA), Institute for Astronomy, University of Edinburgh, Blackford Hill, Edinburgh, EH9 3HJ, UK 2Institute for Physics and Astronomy, University of Potsdam, 14476 Potsdam, Germany 3Instituto de Astrofísica de Andalucía, IAA-CSIC. c/ Glorieta de la Astronomía s/n, 18008 Granada, Spain 4Department of Astronomy, Nanjing University, Nanjing, 210093 Jiangsu, China 5Department of Astronomy, University of Illinois, 1002 West Green Street, Urbana, IL 61801, USA 6UK Astronomy Technology Centre, Royal Observatory Edinburgh, Blackford Hill, Edinburgh, EH9 3HJ, UK 7Department of Astronomy, University of Wisconsin-Madison, 5534 Sterling, 475 North Charter St., Madison, WI 53706, USA 8Escuela Superior de Física y Matemáticas, IPN, U.P. Adolfo López Mateos, C.P. 07738 D.F., México

We report on a new Be/X-ray pulsar binary located in the Wing of the Small Magellanic Cloud (SMC). The strong pulsed X-ray source was discovered with the Chandra and XMM-Newton X-ray observatories. The X-ray pulse period of 1062 s is consistently determined from both Chandra and XMM-Newton observations, revealing one of the slowest rotating X-ray pulsars known in the SMC. The optical counterpart of the X-ray source is the emission-line star 2dFS 3831. Its B0–0.5(III)e+ spectral type is determined from VLT-FLAMES and 2dF optical spectroscopy, establishing the system as a Be/X-ray binary (Be-XRB). The hard X-ray spectrum is well fitted by a power-law with additional thermal and blackbody components, the latter reminiscent of persistent Be-XRBs. This system is the first evidence of a recent supernova in the low density surroundings of NGC 602. We detect a shell nebula around 2dFS 3831 in Hα and [O iii] images and conclude that it is most likely a supernova remnant. If it is linked to the supernova explosion that created this new X-ray pulsar, its kinematic age of (2–4)×10^{4} yr provides a constraint on the age of the pulsar.

Accepted for publication in MNRAS Letters Available from arXiv:1110.6404

The VLT-FLAMES survey of massive stars: NGC 346-013 as a test case for massive close binary evolution

B.W. Ritchie^{1}, V.E. Stroud^{1,2,3}, C.J. Evans^{4}, J.S. Clark^{1}, I. Hunter^{5}, D.J. Lennon^{6}, N. Langer^{7} and S.J. Smartt^{5}

1Department of Physics and Astronomy, The Open University, Walton Hall, Milton Keynes MK7 6AA, UK 2Faulkes Telescope Project, School of Physics and Astronomy, Cardiff University, Cardiff, CF24 3AA, UK 3Division of Earth, Space and Environment, University of Glamorgan, Pontypridd, CF37 1DL, UK 4UK Astronomy Technology Centre, Royal Observatory Edinburgh, Blackford Hill, Edinburgh, EH9 3HJ, UK 5Department of Physics & Astronomy, Queen’s University Belfast, Belfast BT7 1NN , Northern Ireland, UK 6ESA/STScI, 3700 San Martin Drive, Baltimore, MD 21218, USA 7Argelander-Institut für Astronomie der Universität Bonn, Auf dem Hügel 71, 53121 Bonn, Germany

NGC 346-013 is a peculiar double-lined eclipsing binary in the Small Magellanic Cloud discovered by the VLT-FLAMES survey of massive stars. Spectra obtained with VLT-FLAMES are used to construct a radial velocity curve and photometry obtained with the Faulkes Telescope South is then used to derive orbital parameters, while spectra of the secondary are compared with synthetic spectra from tlusty model atmospheres. The orbital period is found to be 4.20381(12) days, with masses of 19.1 ± 1.0 and 11.9 ± 0.6 M_{⊙}. The primary is a rapidly rotating late-O dwarf while the secondary, an early-B giant, displays near-synchronous rotation and has filled its Roche lobe, implying that it was originally the more massive component with recent mass transfer ‘spinning up’ the primary to nearcritical rotation. Comparison with synthetic spectra finds temperatures of 34.5 kK and 24.5 kK for the primary and secondary respectively, with the nitrogen abundance of the secondary enhanced compared to baseline values for the SMC, consistent with the predictions of models of interacting binaries. NGC 346-013 likely evolved via non-conservative mass transfer in a system with initial masses ∼ 22+15 M_{⊙}, with the well-constrained orbital solution and atmospheric parameters making it an excellent candidate for tailored modelling with binary evolution codes. This system will form a cornerstone in constraining the physics of thermal timescale mass transfer, and the associated mass transfer efficiency, in massive close binary systems.

Accepted for publication in A&A Available from arXiv:1110.6325

Discovery of a giant, highly-collimated jet from Sanduleak’s star in the Large Magellanic Cloud

R. Angeloni^{1}, F. Di Mille^{2}, J. Bland-Hawthorn^{3} and D. Osip^{4}

1Departamento de Astronomía y Astrofísica, Pontificia Universidad Católica de Chile, Chile 2Australian Astronomical Observatory – Carnegie Observatories, Chile 3Sydney Institute for Astronomy, School of Physics, University of Sydney, Australia 4Las Campanas Observatory, Carnegie Observatories, Chile

Highly-collimated gas ejections are among the most dramatic structures in the Universe, observed to emerge from very different astrophysical systems – from active galactic nuclei down to young brown dwarf stars. Even with the huge span in spatial scales, there is convincing evidence that the physics at the origin of the phenomenon, namely the acceleration and collimation mechanisms, is the same in all classes of jets. Here we report on the discovery of a giant, highly-collimated jet from Sanduleak’s star in the Large Magellanic Cloud (LMC). With a physical extent of 14 parsecs at the distance of the LMC, it represents the largest stellar jet ever discovered, and the first resolved stellar jet beyond the Milky Way. The kinematics and extreme chemical composition of the ejecta from Sanduleak’s star bear strong resemblance with the low-velocity remnants of SN 1987A and with the outer filaments of the most famous supernova progenitor candidate, i.e., η Carinae. Moreover, the precise knowledge of the jet’s distance implies that it will be possible to derive accurate estimates of most of its physical properties. Sanduleak’s bipolar outflow will thus become a crucial test-bed for future theoretical modeling of astrophysical jets.

Accepted for publication in Astrophysical Journal Letters Available from arXiv:1111.0848

Surveying the Agents of Galaxy Evolution in the tidally stripped, low metallicity Small Magellanic Cloud (SAGE-SMC). I. Overview

Karl Gordon^{1} and SAGE-SMC Legacy Team

1Space Telescope Science Institute, USA

The Small Magellanic Cloud (SMC) provides a unique laboratory for the study of the lifecycle of dust given its low metallicity (∼ 1/5 solar) and relative proximity (∼ 60 kpc). This motivated the SAGE-SMC (Surveying the Agents of Galaxy Evolution in the tidally stripped, low metallicity Small Magellanic Cloud) Spitzer Legacy program with the specific goals of studying the amount and type of dust in the present interstellar medium, the sources of dust in the winds of evolved stars, and how much dust is consumed in star formation. This program mapped the full SMC (30 deg^{2}) including the Body, Wing, and Tail in seven bands from 3.6 to 160 μm using IRAC and MIPS on the Spitzer Space Telescope. The data were reduced and mosaicked, and the point sources were measured using customized routines specific for large surveys. We have made the resulting mosaics and point-source catalogs available to the community. The infrared colors of the SMC are compared to those of other nearby galaxies and the 8 μm/24 μm ratio is somewhat lower than the average and the 70 μm/160 μm ratio is somewhat higher than the average. The global infrared spectral energy distribution (SED) shows that the SMC has approximately 1/3 the aromatic emission/polycyclic aromatic hydrocarbon abundance of most nearby galaxies. Infrared color–magnitude diagrams are given illustrating the distribution of different asymptotic giant branch stars and the locations of young stellar objects. Finally, the average SED of H ii/star formation regions is compared to the equivalent Large Magellanic Cloud average H ii/star formation region SED. These preliminary results will be expanded in detail in subsequent papers.

Published in AJ, 142, 102 (2011) Available from arXiv:1107.4313 and from http://adsabs.harvard.edu/abs/2011AJ....142..102G

New 20-cm radio-continuum study of the Small Magelanic Cloud: Part I – Images

G.F. Wong^{1}, M.D. Filipović^{1}, E.J. Crawford^{1}, A. De Horta^{1}, T. Galvin^{1}, D. Drašković,^{1} and J.L. Payne^{1}

1University of Western Sydney, Australia

We present and discuss new high-sensitivity and resolution radiocontinuum images of the Small Magellanic Cloud (SMC) at λ = 20 cm (ν = 1.4 GHz). The new images were created by merging 20-cm radio-continuum archival data, from the Australian Telescope Compact Array and the Parkes radio-telescope. Our images span from ∼ 10′′ to ∼ 150′′ in resolution and sensitivity of r.m.s. ∼ 0.5 mJy beam^{−1}. These images will be used in future studies of the SMC’s intrinsic sources and its overall extended structure.

Published in Serbian Astronomical Journal Available from arXiv:1105.3257 and from http://adsabs.harvard.edu/abs/2011SerAJ.182...43W

New 20-cm radio-continuum study of the Small Magellanic Cloud: Part II – Point sources

G.F. Wong^{1}, M.D. Filipović^{1}, E.J. Crawford^{1}, N.F.H. Tothill^{1}, A.Y. De Horta^{1}, D. Drašković^{1}, T.J. Galvin^{1}, J.D. Collier^{1} and J.L. Payne^{1}

1University of Western Sydney, Australia

We present a new catalogue of radio-continuum sources in the field of the Small Magellanic Cloud (SMC). This catalogue contains sources previously not found in 2370 MHz (λ = 13 cm) with sources found at 1400 MHz (λ = 20 cm) and 843 MHz (λ = 36 cm). 45 sources have been detected at 13 cm, with 1560 sources at 20 cm created from new high sensitivity and resolution radio-continuum images of the SMC at 20 cm from Paper I. We also created a 36 cm catalogue to which we listed 1689 radio-continuum sources.

Accepted for publication in Serbian Astronomical Journal Available from arXiv:1111.0160

New 6 and 3-cm radio-continuum maps of the Small Magellanic Cloud. Part I – The maps

E.J. Crawford^{1}, M.D. Filipović^{1}, A.Y. De Horta^{1}, G.F. Wong^{1}, N.F.H. Tothill^{1}, D. Drašković^{1}, J.D. Collier^{1} and T.J. Galvin^{1}

1University of Western Sydney, Australia

We present new 6 and 3-cm radio-continuum maps of the Small Magellanic Cloud (SMC), created with the ”peeling” technique and a joint deconvolution. The maps have resolutions of 30′′ and 20′′ and r.m.s. noise of 0.7 and 0.8 mJy beam^{−1} at 6 and 3 cm, respectively. These maps will be used for future studies of the SMC’s radio source population and overall extended structure.

Accepted for publication in Serbian Astronomical Journal Available from arXiv:1110.5096

Is there a mass discrepancy in the Cepheid binary OGLE-LMC-CEP 0227?

Hilding R. Neilson^{1} and Norbert Langer^{1}

1Argelander Institute for Astronomy, Germany

Context: The Cepheid mass discrepancy, the difference between masses predicted from stellar evolution and stellar pulsation calculations, is a challenge for the understanding of stellar astrophysics. Recent models of the eclipsing binary Cepheid OGLE-LMC-CEP 0227 have suggested that the discrepancy may be resolved. Aims: We explore for what physical parameters do stellar evolution models agree with the measured properties of OGLE-LMC-CEP 0227 and compare to canonical stellar evolution models assuming no convective core overshooting. Methods: We construct state-of-the-art stellar evolution models for varying mass, metallicity, and convective core overshooting and compare the stellar evolution predictions with the observed properties. Results: The observed mass, effective temperature, and radius of the two stars in the binary system are well fit by numerous combinations of physical parameters, suggesting a Cepheid mass discrepancy of 10–20% relative to canonical stellar evolution models. Conclusions: The properties of the observed binary Cepheid suggest that the Cepheid mass discrepancy is still a challenge and requires more specific observations, such as the rate of period change, to better constrain and understand the necessary physics for stellar evolution models to resolve the discrepancy.

Accepted for publication in Astronomy & Astrophysics Available from arXiv:1110.6657

Photometric determination of the mass accretion rates of pre-main sequence stars. III. Results in the Large Magellanic Cloud

L.Spezzi^{1}, G. De Marchi^{1}, N. Panagia^{2}, A. Sicilia-Aguilar^{3} and B. Ercolano^{4}

1European Space Agency (ESTEC), P.O. Box 299, 2200 AG Noordwijk, The Netherlands 2Space Telescope Science Institute, 3700 San Martin Drive, Baltimore, MD 21218, USA 3Departamento de Física Teórica, Universidad Autónoma de Madrid, Cantoblanco 28049, Madrid, Spain 4Ludwig-Maximilians-Universität, University Observatory Munich, Scheinerstr. 1, D-81679 Munchen, Germany

We present a multi-wavelength study of three star forming regions, spanning the age range 1–14 Myr, located between the 30 Doradus complex and supernova SN 1987A in the Large Magellanic Cloud (LMC). We reliably identify about 1000 pre-main sequence (PMS) star candidates actively undergoing mass accretion and estimate their stellar properties and mass accretion rate (Ṁ). Our measurements represent the largest Ṁ dataset of low-metallicity stars presented so far. As such, they offer a unique opportunity to study on a statistical basis the mass accretion process in the LMC and, more in general, the evolution of the mass accretion process around low-metallicity stars. We find that the typical Ṁ of PMS stars in the LMC is higher than for galactic PMS stars of the same mass, independently of their age. Taking into account the caveats of isochronal age and Ṁ estimates, the difference in Ṁ between the LMC and our Galaxy appears to be about an order of magnitude. We review the main mechanisms of disk dispersal and find indications that typically higher Ṁ are to be expected in low-metallicity environments. However, many issues of this scenario need to be clarified by future observations and modeling. We also find that, in the mass range 1–2 M_{⊙}, the Ṁ of PMS stars in the LMC increases with stellar mass as Ṁ ∝ M_{⋆}^{b} with b ≈ 1, i.e. slower than the second power low found for galactic PMS stars in the same mass regime.

Accepted for publication in Monthly Notices of the Royal Astronomical Society Available from arXiv:1111.0835

Runaway massive stars from R 136: VFTS 682 is very likely a “slow runaway”

Sambaran Banerjee^{1}, Pavel Kroupa^{1} and Seungkyung Oh^{1}

1Argelander-Institut für Astronomie, Auf dem Hügel 71, D-53121, Bonn, Germany

We conduct a theoretical study on the ejection of runaway massive stars from R 136 – the central massive, star-burst cluster in the 30 Doradus complex of the Large Magellanic Cloud. Specifically, we investigate the possibility of the very massive star (VMS) VFTS 682 being a runaway member of R 136. Recent observations of the above VMS, by virtue of its isolated location and its moderate peculiar motion, have raised the fundamental question whether isolated massive star formation is indeed possible. We perform the first realistic N-body computations of fully mass-segregated R 136-type star clusters in which all the massive stars are in primordial binary systems. These calculations confirm that the dynamical ejection of a VMS from a R 136-like cluster, with kinematic properties similar to those of VFTS 682, is common. Hence the conjecture of isolated massive star formation is unnecessary to account for this VMS. Our results are also quite consistent with the ejection of 30 Dor 016, another suspected runaway VMS from R 136. We further note that during the clusters’ evolution, mergers of massive binaries produce a few single stars per cluster with masses significantly exceeding the canonical upper-limit of 150 M_{⊙}. The observations of such single super-canonical stars in R 136, therefore, do not imply an IMF with an upper limit greatly exceeding the accepted canonical 150 M_{⊙} limit, as has been suggested recently, and they are consistent with the canonical upper limit.

Accepted for publication in The Astrophysical Journal Available from arXiv:1111.0291

Nitrogen line spectroscopy in O-stars – II. Surface nitrogen abundances for O-stars in the Large Magellanic Cloud

J.G. Rivero González^{1}, J. Puls^{1}, F. Najarro^{2} and I. Brott^{3}

1Universitätssternwarte München, Scheinerstr. 1, 81679 München, Germany 2Centro de Astrobiología (CSIC-INTA), Ctra. Torrejón a Ajalvir km 4, 28850 Torrejón de Ardoz, Spain 3University of Vienna, Department of Astronomy, Türkenschanzstr. 17, 1180 Vienna, Austria

Context: Nitrogen is a key element to test the impact of rotational mixing on evolutionary models of massive stars. Recent studies of the nitrogen surface abundance in B-type stars within the VLT-FLAMES survey of massive stars have challenged part of the corresponding predictions. To obtain a more complete picture of massive star evolution, and to allow for further constraints, these studies need to be extended to O-stars. Aims: This is the second paper in a series aiming at the analysis of nitrogen abundances in O-type stars, to enable further constraints on the early evolution of massive stars. In this paper, we investigate the N iv λ4058 emission line formation, provide nitrogen abundances for a substantial O-star sample in the Large Magellanic Cloud, and compare our (preliminary) findings with recent predictions from stellar evolutionary models. Methods: Stellar and wind parameters of our sample stars are determined by line profile fitting of hydrogen, helium and nitrogen lines, exploiting the corresponding ionization equilibria. Synthetic spectra are calculated by means of the NLTE atmosphere/spectrum synthesis code fastwind, using a new nitrogen model atom. We derive nitrogen abundances for 20 O- and 5 B-stars, by analyzing all nitrogen lines (from different ionization stages) present in the available optical spectra. Results: The dominating process responsible for emission at N iv λ4058 in O-stars is the strong depopulation of the lower level of the transition, which increases as a function of Ṁ. Unlike the N iii triplet emission, resonance lines do not play a role for typical mass-loss rates and below. We find (almost) no problem in fitting the nitrogen lines, in particular the ‘f’ features. Only for some objects, where lines from N iii/N iv/N v are visible in parallel, we need to opt for a compromise solution. For five objects in the early B-/late O-star domain which have been previously analyzed by different methods and model atmospheres, we derive consistent nitrogen abundances. The bulk of our sample O-stars seems to be strongly nitrogen-enriched, and a clear correlation of nitrogen and helium enrichment is found. By comparing the nitrogen abundances as a function of vsini (’Hunter-plot’) with tailored evolutionary calculations, we identify a considerable number of highly enriched objects at low rotation. Conclusions: Our findings seem to support the basic outcome of previous B-star studies within the VLT-FLAMES survey. Due to the low initial abundance, the detection of strong nitrogen enrichment in the bulk of O-stars indicates that efficient mixing takes place already during the very early phases of stellar evolution of LMC O-stars. For tighter constraints, however, upcoming results from the VLT-FLAMES Tarantula survey need to be waited for, comprising a much larger number of O-stars that will be analyzed based on similar methods as presented here.

Accepted for publication in Astronomy & Astrophysics Available from arXiv:1110.5148 and from http://www.usm.uni-muenchen.de/people/puls/electronic prints.html

The VMC survey IV. The LMC star formation history and disk geometry from four VMC tiles

S. Rubele^{1,2}, L. Kerber^{3}, L. Girardi^{1}, M.-R. Cioni^{4,5}, P. Marigo^{1}, S. Zaggia^{1}, K. Bekki^{6}, R. de Grijs^{7,8}, J. Emerson^{9}, M.A.T. Groenewegen^{10}, M. Gullieuszik^{10}, V. Ivanov^{11}, B. Miszalski^{12,13}, J.M. Oliveira^{14}, B. Tatton^{14}

and J.Th. van Loon^{14}

1Osservatorio Astronomico di Padova – INAF, Vicolo dell’Osservatorio 5, 35122 Padova, Italy 2Dipartimento di Astronomia, Università di Padova, Vicolo dell’Osservatorio 2, 35122 Padova, Italy 3Universidade Estadual de Santa Cruz, Rodovia Ilhéus-Itabuna, km. 16 – 45662-000 Ilhéus, Bahia, Brazil 4University of Hertfordshire, Physics Astronomy and Mathematics, Hatfield AL10 9AB, UK 5University Observatory Munich, Scheinerstraße 1, 81679 München, Germany 6ICRAR M468, The University of Western Australia, 35 Stirling Highway, Crawley, WA 6009, Australia 7Kavli Institute for Astronomy and Astrophysics, Peking University, Yi He Yuan Lu 5, Hai Dian District, 100871 Beijing, China 8Department of Astronomy and Space Science, Kyung Hee University, Yongin-shi, 449-701 Kyungki-do, South Korea 9Astronomy Unit, Queen Mary University of London, Mile End Road, London E1 4NS, UK 10Royal Observatory of Belgium, Ringlaan 3, 1180 Brussels, Belgium 11European Southern Observatory, Av. Alonso de Córdoba 3107, Casilla 19, Santiago, Chile 12South African Astronomical Observatory, P.O. Box 9, 7935 Observatory, South Africa 13Southern African Large Telescope Foundation, P.O. Box 9, 7935 Observatory, South Africa 14Lennard-Jones Laboratories, Keele University, ST5 5BG, UK

We derive the star formation history (SFH) for several regions of the Large Magellanic Cloud (LMC), using deep near-infrared data from the VISTA near-infrared Y JK_{s} survey of the Magellanic system (VMC). The regions include three almost-complete 1.4 deg^{2} tiles located ∼ 3.5^{∘} away from the LMC centre in distinct directions. They are split into 21.′0 × 21.′5 (0.12 deg^{2} ) subregions, and each of these is analysed independently. To this dataset, we add two 11.′3 × 11.′3 (0.036 deg^{2} ) subregions selected based on their small and uniform extinction inside the 30 Doradus tile. The SFH is derived from the simultaneous reconstruction of two different colour–magnitude diagrams (CMDs), using the minimization code StarFISH together with a database of ”partial models” representing the CMDs of LMC populations of various ages and metallicities, plus a partial model for the CMD of the Milky Way foreground. The distance modulus (m−M)_{0} and extinction A_{V} is varied within intervals ∼ 0.2 and ∼ 0.5 mag wide, respectively, within which we identify the best-fitting star formation rate SFR(t) as a function of lookback time t, age–metallicity relation (AMR), (m − M)_{0} and A_{V}. Our results demonstrate that VMC data, due to the combination of depth and little sensitivity to differential reddening, allow the derivation of the space-resolved SFH of the LMC with unprecedented quality compared to previous wide-area surveys. In particular, the data clearly reveal the presence of peaks in the SFR(t) at ages log(t/yr) ≃ 9.3 and 9.7, which appear in most of the subregions. The most recent SFR(t) is found to vary greatly from subregion to subregion, with the general trend of being more intense in the innermost LMC, except for the tile next to the N 11 complex. In the bar region, the SFR(t) seems remarkably constant over the time interval from log(t/yr) ≃ 8.4 to 9.7. The AMRs, instead, turn out to be remarkably similar across the LMC. Thanks to the accuracy in determining the distance modulus for every subregion – with typical errors of just ∼ 0.03 mag – we make a first attempt to derive a spatial model of the LMC disk. The fields studied so far are fit extremely well by a single disk of inclination i = 26.^{∘}2 ± 2.^{∘}0, position angle of the line of nodes θ_{0} = 129.^{∘}1 ± 13.^{∘}0, and distance modulus of (m − M)_{0} = 18.470 ± 0.006 mag (random errors only) up to the LMC centre. We show that once the (m − M)_{0} values or each subregion are assumed to be identical to those derived from this best-fitting plane, systematic errors in the SFR(t) and AMR are reduced by a factor of about two.

Accepted for publication in Astronomy & Astrophysics Available from arXiv:1110.5852

The VLT-FLAMES Tarantula Survey: The fastest rotating O-type star and shortest period LMC pulsar – remnants of a supernova disrupted

binary?

P.L. Dufton^{1}, P.R. Dunstall^{1}, C.J. Evans^{2}, I. Brott^{3}, M. Cantiello^{4,5}, A. de Koter^{6}, S.E. de Mink^{7}, M. Fraser^{1}, V. Hénault-Brunet^{8}, I.D. Howarth^{9}, N. Langer^{4}, D.J. Lennon^{10}, N. Markova^{11}, H. Sana^{6} and W.D. Taylor^{8}

1Astrophysics Research Centre, School of Mathematics and Physics, Queen’s University Belfast, Belfast BT7 1NN, Northern Ireland, UK 2UK Astronomy Technology Centre, Royal Observatory Edinburgh, Blackford Hill, Edinburgh, EH9 3HJ, UK 3University of Vienna, Department of Astronomy, Türkenschanzstr. 17, A-1180 Vienna, Austria 4Argelander Institut für Astronomie der Universität Bonn, Auf dem Hügel 71, 53121 Bonn, Germany 5Kavli Institute for Theoretical Physics, Kohn Hall, University of California, Santa Barbara, CA 93106, USA 6Astronomical Institute ‘Anton Pannekoek’, University of Amsterdam, Postbus 94249, 1090 GE, Amsterdam, The Netherlands 7Space Telescope Science Institute, 3700 San MartinDrive, Baltimore, MD 21218, USA 8Scottish Universities Physics Alliance, Institute for Astronomy, University of Edinburgh, Royal Observatory Edinburgh, Blackford Hill, Edinburgh, EH9 3HJ, UK 9Department of Physics & Astronomy, University College London, Gower Street, London, WC1E 6BT, UK 10ESA, Space Telescope Science Institute, 3700 San MartinDrive, Baltimore, MD 21218, USA 11Institute of Astronomy with NAO, Bulgarian Academy of Sciences, P.O. Box 136, 4700 Smoljan, Bulgaria

We present a spectroscopic analysis of an extremely rapidly rotating late O-type star, VFTS 102, observed during a spectroscopic survey of 30 Doradus. VFTS 102 has a projected rotational velocity larger than 500 km s^{−1} and probably as large as 600 km s^{−1}; as such it would appear to be the most rapidly rotating massive star currently identified. Its radial velocity differs by 40 km s^{−1} from the mean for 30 Doradus, suggesting that it is a runaway. VFTS 102 lies 12 pc from the X-ray pulsar PSR J0537−6910 in the tail of its X-ray diffuse emission. We suggest that these objects originated from a binary system with the rotational and radial velocities of VFTS 102 resulting from mass transfer from the progenitor of PSR J0537−691 and the supernova explosion respectively.

Accepted for publication in Astrophysical Journal Letters Available from arXiv:1111.0157

Evolutionary models of the optical component of the LMC X-1/Star 32 binary system

Janusz Ziółkowski ^{1}

1Copernicus Astronomical Center, Warsaw, Poland

Calculations carried out to model the evolution of Star 32 under different assumptions about the stellar wind mass-loss rate provide robust limits on the present mass of the star. The obtained range is 31 to 35.5 M_{⊙}, which is in very good agreement with the orbital solution of Orosz et al., namely 28.3 to 35.3 M_{⊙}. The initial mass of Star 32 had to be in the range 35 to 40 M_{⊙} and the present age of the system is 3.7 to 4.0 Myr.

Accepted for publication in MNRAS Available from arXiv:1110.5312

Theoretical Cepheid period–luminosity and period–color relations in Spitzer IRAC bands

Chow-Choong Ngeow^{1}, Marcella Marconi^{2}, Ilaria Musella^{2}, Michele Cignoni^{3} and Shashi M. Kanbur^{4}

1NCU, Taiwan 2Osservatorio Astronomico di Capodimonte, Italy 3Bologna University, Italy 4SUNY-Oswego, USA

In this paper the synthetic period–luminosity (P–L) relations in Spitzer’s IRAC bands, based on a series of theoretical pulsation models with varying metal and helium abundance, were investigated. Selected sets of these synthetic P–L relations were compared to the empirical IRAC band P–L relations recently determined from Galactic and Magellanic Clouds Cepheids. For the Galactic case, synthetic P–L relations from model sets with (Y = 0.26, Z = 0.01), (Y = 0.26, Z = 0.02) and (Y = 0.28, Z = 0.02) agree with the empirical Galactic P–L relations derived from the Hubble Space Telescope parallaxes. For Magellanic Cloud Cepheids, the synthetic P–L relations from model sets with (Y = 0.25, Z = 0.008) agree with both of the empirical Large Magellanic Cloud (LMC) and Small Magellanic Cloud (SMC) P–L relations. Analysis of the synthetic P–L relations from all model sets suggested that the IRAC band P–L relations may not be independent of metallicity, as the P–L slopes and intercepts could be affected by the metallicity and/or helium abundance. We also derive the synthetic period-color (P–C) relations in the IRAC bands. Non-vanishing synthetic P–C relations were found for certain combinations of IRAC band filters and metallicity. However, the synthetic P–C relations disagreed with the [3.6]–[8.0] P–C relation recently found for the Galactic Cepheids. The synthetic [3.6]–[4.5] P–C slope from (Y = 0.25, Z = 0.008) model set, on the other hand, is in excellent agreement to the empirical LMC P–C counterpart, if a period range of 1.0 < log(P) < 1.8 is adopted.

Accepted for publication in Astrophysical Journal Available from arXiv:1111.1791

HST-COS observations of hydrogen, helium, carbon and nitrogen emission from the SN 1987A reverse shock

Kevin France^{1} and et al.

1University of Colorado, USA

We present the most sensitive ultraviolet observations of Supernova 1987A to date. Imaging spectroscopy from the Hubble Space Telescope-Cosmic Origins Spectrograph shows many narrow (Δv ∼ 300 km s^{−1}) emission lines from the circumstellar ring, broad (Δv ∼ 10–20 × 10^{3} km s^{−1}) emission lines from the reverse shock, and ultraviolet continuum emission. The high signal-to-noise (> 40 per resolution element) broad Lyα emission is excited by soft X-ray and EUV heating of mostly neutral gas in the circumstellar ring and outer supernova debris. The ultraviolet continuum at λ > 1350 Å can be explained by H i 2-photon emission from the same region. We confirm our earlier, tentative detection of N v λ1240 emission from the reverse shock and we present the first detections of broad He ii λ1640, C iv λ1550, and N iv] λ1486 emission lines from the reverse shock. The helium abundance in the high-velocity material is He/H = 0.14 ± 0.06. The N v/Hα line ratio requires partial ion–electron equilibration (T_{e}/T_{p} ∼ 0.14–0.35). We find that the N/C abundance ratio in the gas crossing the reverse shock is significantly higher than that in the circumstellar ring, a result that may be attributed to chemical stratification in the outer envelope of the supernova progenitor. The N/C abundance ratio may have been stratified prior to the ring expulsion, or this result may indicate continued CNO processing in the progenitor subsequent to the expulsion of the circumstellar ring.

Accepted for publication in ApJ Available from arXiv:1111.1735

The rich circumstellar chemistry of SMP LMC 11

Sarah Malek^{1}, Jan Cami^{1,2} and Jerónimo Bernard-Salas^{3}

1University of Western Ontario, London, Canada 2SETI Institute, Mountain View, USA 3Institut d’Astrophysique Spatiale, CNRS/Université Paris-Sud, France

Carbon-rich evolved stars from the asymptotic giant branch to the planetary nebula phase are characterized by a rich and complex carbon chemistry in their circumstellar envelopes. A peculiar object is the preplanetary nebula SMP LMC 11, whose Spitzer-IRS spectrum shows remarkable and diverse molecular absorption bands. To study how the molecular composition in this object compares to our current understanding of circumstellar carbon chemistry, we modeled this molecular absorption. We find high abundances for a number of molecules, perhaps most notably benzene. We also confirm the presence of propyne (CH_{3}C_{2}H) in this spectrum. Of all the cyanopolyynes, only HC_{3}N is evident; we can detect at best a marginal presence of HCN. From comparisons to various chemical models, we can conclude that SMP LMC 11 must have an unusual circumstellar environment (a torus rather than an outflow).

Accepted for publication in The Astrophysical Journal Available from arXiv:1111.2533

Evolution and nucleosynthesis of AGB stars in three Magellanic Cloud clusters

D. Kamath^{1}, A.I. Karakas^{1} and P.R. Wood^{1}

1Research School of Astronomy & Astrophysics, Mount Stromlo Observatory, Weston Creek ACT 2611, Australia

We present stellar evolutionary sequences for asymptotic giant branch (AGB) stars in the Magellanic Cloud clusters NGC 1978, NGC 1846 and NGC 419. The new stellar models for the three clusters match the observed effective temperatures on the giant branches, the oxygen-rich to carbon-rich transition luminosities, and the AGB-tip luminosities. A major finding is that a large amount of convective overshoot (up to 3 pressure scale heights) is required at the base of the convective envelope during third dredge-up in order to get the correct oxygen-rich to carbon-rich transition luminosity. The stellar evolution sequences are used as input for detailed nucleosynthesis calculations. For NGC 1978 and NGC 1846 we compare our model results to the observationally derived abundances of carbon and oxygen. We find that additional mixing processes (extra-mixing) are required to explain the observed abundance patterns. For NGC 1846 we conclude that non-convective extra-mixing processes are required on both the RGB and the AGB, in agreement with previous studies. For NGC 1978 it is possible to explain the C/O and ^{12}C/^{13}C abundances of both the O-rich and the C-rich AGB stars by assuming that the material in the intershell region contains high abundances of both C and O. This may occur during a thermal pulse when convective overshoot at the inner edge of the flash-driven convective pocket dredges C and O from the core to the intershell. For NGC 419 we provide our predicted model abundance values although there are currently no published observed abundance studies for the AGB stars in this cluster.

Accepted for publication in ApJ Available from arXiv:1111.1722

Interstellar H i and H_{2} in the Magellanic Clouds: An expanded sample based on UV absorption-line data

Daniel E. Welty^{1}, Rui Xue^{2} and Tony Wong^{2}

1University of Chicago, USA 2University of Illinois at Urbana/Champaign, USA

We have determined column densities of H i and/or H_{2} for sight lines in the Magellanic Clouds from archival HST and FUSE spectra of H i Lyman-α and H_{2} Lyman-band absorption. Together with some similar data from the literature, we now have absorption-based N(H i) and/or N(H_{2}) for 285 LMC and SMC sight lines (114 with a detection or limit for both species) – enabling more extensive, direct, and accurate determinations of molecular fractions, gas-to-dust ratios, and elemental depletions in these two nearby, low-metallicity galaxies. For sight lines where the N(H i) estimated from 21 cm emission is significantly higher than the value derived from Lyman-α absorption (presumably due to emission from gas beyond the target stars), integration of the 21 cm profile only over the velocity range seen in Na i or H_{2} absorption generally yields much better agreement. Conversely, N(21 cm) can be lower than N(Lyα) by factors of 2–3 in some LMC sight lines – suggestive of small-scale structure within the 21 cm beam(s) and/or some saturation in the emission. The mean gas-to-dust ratios obtained from N(H_{tot})/E(B − V) are larger than in our Galaxy, by factors of 2.8–2.9 in the LMC and 4.1–5.2 in the SMC – i.e., factors similar to the differences in metallicity. The N(H_{2})/E(B − V) ratios are more similar in the three galaxies, but with considerable scatter within each galaxy. These data may be used to test models of the atomic-to-molecular transition at low metallicities and predictions of N(H_{2}) based on comparisons of 21 cm emission and the IR emission from dust.

Accepted for publication in ApJ Available from arXiv:1111.3674

Period–luminosity relations for Small Magellanic Cloud Cepheids based on AKARI archival data

Chow-Choong Ngeow^{1}, Danielle M. Citro^{2} and Shashi M. Kanbur^{2}

1NCU, Taiwan 2SUNY-Oswego, USA

In this work we matched the AKARI archival data to the Optical Gravitational Lensing Experiment-III (OGLE-III) catalog to derive the mid-infrared period–luminosity (P–L) relations for Small Magellanic Cloud (SMC) Cepheids. Mismatched AKARI sources were eliminated using random-phase colors obtained from the full I-band light curves from OGLE-III. It was possible to derive P–L relations in the N3 and N4 bands only, although the S7, S11, L15, and L24 band data were also tested. Random-phase correction was included when deriving the P–L relation in the N3 and N4 bands using the available time of observations from AKARI data. The final adopted P–L relations were: N3 = −3.370 log P + 16.527 and N4 = −3.402 log P + 16.556. However, these P–L relations may be biased due to the small number of Cepheids in the sample.

Accepted for publication in MNRAS Available from arXiv:1111.4347

Conference Papers

The IMF of field OB stars in the Small Magellanic Cloud

J.B. Lamb^{1}, M.S. Oey^{1}, A.S. Graus^{1} and D.M. Segura-Cox^{1}

1University of Michigan Department of Astronomy, 830 Dennison Bldg., Ann Arbor, MI 48109-1042, USA

The population of field OB stars are an important component of a galaxy’s stellar content, representing 20–30% of the massive stars. To study this population, we have undertaken the Runaways and Isolated O Type Star Spectroscopic Survey of the SMC (RIOTS4). RIOTS4 surveys a spatially complete sample of > 350 field OB stars in the Small Magellanic Cloud and will serve as a key probe of runaways, binaries, and the stellar IMF in the field massive star population. Here, we focus on the field IMF, which provides an empirical probe of the star-forming process and is a fundamental property of a stellar population. Together with photometry from the OGLE survey, RIOTS4 will yield a definitive stellar IMF for the SMC field massive star population. We present preliminary results that suggests the field IMF is much steeper, Γ = 2.9, than the canonical stellar IMF of Γ = 1.35. Despite the steep slope, we see no evidence of a stellar upper mass limit, up to our highest mass star of 65 M_{⊙}.

Oral contribution, published in ”Four Decades of Research on Massive Stars”, eds. L. Drissen, C. Robert, and N. St-Louis, ASP Conference Series Available from arXiv:1109.6655

Discovery of a new Wolf–Rayet star using SAGE-LMC

Vasilii V. Gvaramadze^{1}, André-Nicolas Chené^{2,3}, Alexei Y. Kniazev^{4} and Olivier Schnurr^{5}

1Sternberg Astronomical Institute, Moscow State University, Universitetskij Pr. 13, Moscow 119992, Russia 2Departamento de Astronomía, Universidad de Concepción, Casilla 160-C, Chile 3Universidad de Valparaiso, Departamento de Física y Astronomía, Avenida Gran Bretaña 1111, Valparaiso, Chile, Chile 4South African Astronomical Observatory and Southern African Large Telescope Foundation, P.O. Box 9, 7935 Observatory, Cape Town, South Africa 5AIP, An der Sternwarte 16, 14482 Potsdam, Germany

We report the first-ever discovery of an extragalactic Wolf–Rayet (WR) star with Spitzer. A new WR star in the Large Magellanic Cloud (LMC) was revealed via detection of its circumstellar shell using 24 μm images obtained in the framework of the Spitzer Survey of the Large Magellanic Cloud (SAGE-LMC). Subsequent spectroscopic observations with the Gemini South resolved the central star in two components, one of which is a WN3b+abs star, while the second one is a B0V star. We consider the lopsided brightness distribution over the circumstellar shell as an indication that the WR star is a runaway and use this interpretation to identify a possible parent cluster of the star.

Poster contribution, published in ”Four decades of research on massive stars”, in honor of Tony Moffat, 11–15 July 2011, Saint-Michel-des-Saints, Quebec Available from arXiv:1110.0126

Planetary Nebulae in the VISTA Magellanic Cloud (VMC) survey

B. Miszalski^{1,2}, R. Napiwotzki^{3}, M.-R.L. Cioni^{3,4}, M.A.T. Groenewegen^{5}, J.M. Oliveira^{6}, A. Udalski^{7} and J. Nie^{8,9}

1South African Astronomical Observatory, South Africa 2Southern African Large Telescope Foundation 3University of Hertfordshire, UK 4University Observatory Munich, Germany 5Royal Observatory of Belgium, Belgium 6Keele University, UK 7Warsaw University Observatory, Poland 8Australian National University, Australia 9Beijing Normal University, China

The multi-epoch Y JK_{s} sub-arcsecond photometry of the VMC survey provides a long anticipated deep near-infrared (NIR) window into further understanding the stellar populations of the Magellanic Clouds. The first year of observations consisted of six tiles covering ∼ 9% of the Large Magellanic Cloud (LMC) survey region and contains 102 objects previously classified as planetary nebulae (PNe). A large proportion of the sample were found to be contaminated by non-PNe. These initial results underline the importance of establishing a clean catalogue of LMC PNe before they are applied in areas such as the planetary nebula luminosity function (PNLF) and searches for binary central stars. As the VMC survey progresses it will play a fundamental role in cleaning extant PN catalogues and a complementary role in the discovery of new PNe.

Poster contribution, published in IAUS 283 Available from arXiv:1110.1829

Recent progress of Cepheid research at National Central University: From Spitzer to Kepler

Chow-Choong Ngeow^{1}

1NCU, Taiwan

In this presentation I summarize recent work on Cepheid research carried out at the National Central University. The mid-infrared period–luminosity (P–L) relations for Cepheids are important in the James Webb Space Telescope era for distance scale work, as the relations have potential to derive the Hubble constant within ∼ 2% accuracy – a critical constraint in the precision cosmology. Consequently, we have derived the mid-infrared P–L relations for Cepheids in the Large and Small Magellanic Clouds, using archival data from the Spitzer Space Telescope. Kepler Space Telescope is a NASA mission to search for Earth-size and larger planets around Sun-like stars, by observing continuously the stars in a dedicated patch of the sky. As a result, the almost un-interrupted observation is also used for stellar variability and asteroseismological study. However, Kepler observations are carried out with a single broad-band filter, hence ground-based follow-up observation needed to complement Kepler light curves to fully characterize the properties of the target stars. Here I present the ground-based optical follow-up observations for two Cepheid candidates located within the Kepler’s field-of-view. Together with Kepler light curves, our ground-based data rule out V2279 Cyg being a Cepheid. Hence V1154 Cyg is the only Cepheid in the Kepler’s field.

Oral contribution, published in ”9^{th} Pacific Rim Conference on Stellar Astrophysics” (PRCSA2011), Lijiang, China, April 2011 Available from arXiv:1111.2094

Period-luminosity relations for Magellanic Clouds Cepheids based on OGLE-III data: A comparison

Chow-Choong Ngeow^{1}

1NCU, Taiwan

The period–luminosity (P–L) relation for Cepheid variables is important in modern astrophysics. In this work, we present the multi-band P–L relations derived from the Large Magellanic Cloud (LMC) and Small Magellanic Cloud (SMC) Cepheids, based on the latest release of OGLE-III catalogs. In addition to the V I band mean magnitudes adopted from OGLE-III catalogs, we also cross-matched the LMC and SMC Cepheids to the 2MASS point source catalogs and publicly available Spitzer catalogs from SAGE program. Mean magnitudes for these Cepheids were corrected for extinction using available extinction maps. When comparing the P–L slopes, we found that the P–L slopes in these two galaxies are consistent with each others within ∼ 2.5σ level.

Poster contribution, published in ”11^{th} Asian–Pacific Regional IAU Meeting” (APRIM2011), Chiang-Mai, Thailand, July 2011 Available from arXiv:1111.2095

Statistical tests for the metallicity dependency of the synthetic Cepheid period–luminosity relations in IRAC bands

Chow-Choong Ngeow^{1}, Marcella Marconi^{2}, Ilaria Musella^{2}, Michele Cignoni^{3} and Shashi M. Kanbur^{4}

1NCU, Taiwan 2Osservatorio Astronomico di Capodimonte, Italy 3Bologna University, Italy 4SUNY-Oswego, USA

The mid-infrared (MIR) period–luminosity (P–L) relations for Cepheids will be important in the JWST era, as it holds the promise of deriving the Hubble constant within 2% accuracy. It is expected that the MIR P–L to be insensitive to metallicity. In this work, we test this assumption of metallicity independent of the IRAC band P–L relation by applying well-known statistical methods to the synthetic P–L slopes from a series of pulsating models with known metallicity. The statistical tests suggest that the P–L slopes in MIR are linearly depending on metallicity.

Poster contribution, published in ”11^{th} Asian–Pacific Regional IAU Meeting” (APRIM2011), Chiang-Mai, Thailand, July 2011 Available from arXiv:1111.2096

Synthetic Cepheid period–luminosity & period–color relations in Spitzer’s IRAC bands

Chow-Choong Ngeow^{1}, Marcella Marconi^{2}, Ilaria Musella^{2}, Michele Cignoni^{3}, Massimo Marengo^{4} and Shashi M. Kanbur^{5}

1NCU, Taiwan 2Osservatorio Astronomico di Capodimonte, Italy 3Bologna University, Italy 4Iowa State University, USA 5SUNY-Oswego, USA

The mid-infrared period–luminosity (P–L) relation for Cepheids will be important in the JWST era, as it holds the promise of deriving the Hubble constant within 2% accuracy. We present the synthetic P–L and period–color (P–C) relations, derived from a series of stellar pulsation models with varying helium and metallicity abundance, in the Spitzer IRAC bands. Selected synthetic P–L and P–C relations were compared to the empirical relations derived from Galactic and Large Magellanic Cloud Cepheids. We also present the empirical P–L relations based on the latest distance measurements using infrared surface brightness techniques.

Oral contribution, published in ”20^{th} Stellar Pulsation Conference” (Impact of New Instruments and Insights in Stellar Pulsation), Granada, Spain, September 2011 Available from arXiv:1111.2100

Review Paper

On the apparent lack of Be X-ray binaries with black holes in the Galaxy and in the Magellanic Clouds

Janusz Ziółkowski^{1} and Krzysztof Belczyński^{2}

1Copernicus Astronomical Center, ul. Bartycka 18, 00-716 Warsaw, Poland 2Astronomical Observatory, University of Warsaw, Al. Ujazdowskie 4, 00-478 Warsaw, Poland

In the Galaxy and in the Magellanic Clouds there are 170 Be X-ray binaries known to-date. Out of those, 111 host a neutron star, and for the reminder the nature of a companion is not known. None, so far, is known to host a black hole. This disparity is referred to as a missing Be – black hole X-ray binary problem. The stellar population synthesis calculations following the formation of Be X-ray binaries in the Galaxy (Belczyński and Ziółkowski 2009) demonstrate that there is no problem of the missing Be+BH X-ray binaries for the Galaxy (the expected number of Be – black hole X-ray binaries is 0 to 2, which is entirely consistent with the observed Galactic sample). However, the preliminary calculations for Magellanic Clouds indicate that there is a problem of the missing Be+BH X-ray binaries for the Clouds (the expected number is about 6, while none is observed). We believe, that to remove the discrepancy, one has to take into account a different history of the star formation rate in the Magellanic Clouds, with the respect to the Galaxy. New stellar population synthesis calculations are currently being carried out. An updated (as of November 2011) list of all 170 Be X-ray binaries known presently in the Galaxy and in the Magellanic Clouds is included.

Published in ”Frontier Objects in Astrophysics and Particle Physics” (Vulcano Workshop 2010), eds. F. Giovannelli and G. Mannocchi, Bologna, Italy, 2011 Available from arXiv:1111.2330

Job Adverts

Interstellar Medium Postdoctoral Researcher

Applications are invited for at least one postdoctoral research position at the Space Telescope Science Institute starting as early as March 2012. The successful applicant will work with Drs. Karl Gordon and Julia Roman-Duval on studies of interstellar dust and gas in the Milky Way and nearby galaxies. Potential specific projects include new observations of ultraviolet dust extinction and CO absorption by translucent molecular clouds in the Magellanic Clouds, photometric studies of dust and stars in M 31 as part of the PHAT (Panchromatic Andromeda Hubble Treasury) program, ground-based optical to near-infrared spectroscopic studies of dust extinction in the Milky Way, and infrared studies of dust in the Magellanic Clouds and nearby galaxies using existing Spitzer and Herschel observations (e.g., SAGE-LMC, SAGE-SMC, HERITAGE, SINGS, & KINGFISH program data). This research will involve working with and, if possible, proposing observations with Hubble, Spitzer, Herschel, and ground-based facilities. Independent research in related areas will be supported and encouraged. Research experience in the areas of interstellar dust, molecular clouds, or dusty radiative transfer is desirable.

The position is for two years, with a possible renewal for a third year. STScI, located on Johns Hopkins University Campus in Baltimore, Maryland, offers an excellent benefit package, competitive salaries, and a stimulating work environment. STScI’s pay is commensurate to the year of Ph.D.

Applicants are requested to complete an on-line application through our website at

http://www.stsci.edu/institute/employment

and attach the following materials: curriculum vitae, list of publications, and a brief statement of research interests, accomplishments, and relevant technical expertise in the resume upload section. Please include #11-0084 in the file name. Questions may be directed to kgordon@stsci.edu. Completed applications received by Feb. 1, 2012 are assured of full consideration. Women and members of minority groups are strongly encouraged to apply. EOE/AA/M/F/D/V

Included Benefits:

Comprehensive Medical, Dental, Vision and Prescription Coverage Very generous vacation and leave time, flexible work schedule, casual environment Relocation Assistance Outstanding retirement benefits

See also http://jobregister.aas.org/job view?JobID=40797

Postdoctoral Position in Extragalactic Astrophysics

I’d like to draw your attention to a postdoctoral position we have available at Toledo working with the SPIRE/FTS ”Beyond the Peak” Herschel project and the affiliated KINGFISH Herschel Key Program on nearby galaxies. We are particularly interested in applicants with an IR/radio observational background and/or line modeling expertise. The position can commence as early as Feb., 2012, and further details can be found at the AAS job register:

http://jobregister.aas.org/job view?JobID=40847

The application deadline is 20 Dec.

With warmest regards, Prof. JD Smith University of Toledo

See also http://jobregister.aas.org/node/40847

Announcements

STScI Workshop: Mass Loss Return from Stars to Galaxies

In this small workshop of about 60 participants, we will discuss the topic of mass-loss return to galaxies and the resulting dust and metal enrichment process. The workshop will be about 2.5 days and involve talks and audience-wide discussions. We have plenty of room for contributed talks and posters. As the title suggests, the focus of our workshop is four areas:

1 The parameterizations of mass-loss rates and their basis on facts for both massive stars and intermediate mass stars.

2 The variations in mass loss due to quiescent/smooth, eruptive/episodic, or explosive processes and to the effects of binary companions.

3 How these parameterizations affect both theoretical modeling of stellar evolution and estimates for mass-loss return to galaxies from stellar populations.

4 The composition in dust, metals and total gas of the ejecta and how these are incorporated into dust and chemical evolution of galaxies.

The deadline for registration is March 16^{th}. The deadline for abstract submission is March 2^{nd}.

See also http://www.stsci.edu/institute/conference/stellar-mass-return

Workshop Circumstellar Dynamics at High Resolution

First Announcement Workshop Circumstellar Dynamics at High Resolution Foz do Iguaçu, Brazil, February 27 – March 02, 2012

ESO – Univ. of São Paulo This is the first announcement of a conference on ”Circumstellar Dynamics at High Resolution” to be held February 27 to March 02, 2012. The venue of the meeting will be in Foz do Iguaçu, Brazil.

The Workshop is sponsored by ESO and the University of São Paulo. More detailed information is available on our web page:

http://www.eso.org/sci/meetings/2012/csdyn.html

or by email to csdyninfo@eso.org.

We encourage you to circulate the announcement among your colleagues.

On behalf of the organizers, Alex Carciofi

SCOPE The dynamics of circumstellar (CS) envelopes is an active research frontier that has benefited greatly from the advent of high-resolution observational techniques in the spectral, spatial and temporal domains. The observational discoveries and theoretical results emerging from this field have broad implications for many astrophysical topics, ranging from cosmology (via a better understanding of the progenitors of GRBs, for instance), to star and planet formation (through a better description of CS disk dynamics in which viscosity plays a key role).

The diverse and complex CS environments revealed by these observational techniques are particularly evident near hot high-mass stars, where stellar radiation plays a large if not crucial role in continuously shaping the immediate environment.

High-resolution observations (spatial, spectral, and temporal) have provided important information in several frontline research topics. For example, many hot stars have been shown to be very rapidly rotating, in a regime where geometric deformation and gravity darkening become important. CS structures have not only been resolved spatially, but have been followed over characteristic variation timescales. This dynamical evolution has been modeled for disks and winds: we are now directly observing and measuring the consequence of the physical mechanisms operating within the CS environments. As a result, current observing facilities have allowed the field to progress from a static picture of the CS environment towards understanding its dynamics and concomitant impact on the evolution of the central star.

This workshop aims at bringing together the active community of hot stellar astrophysics, both theoreticians and observers, along the common topic of what can be learned from high resolution observations.

PROGRAM Oral sessions during the meeting will be held on:

1 Circumstellar Disks & Outflows: Theory 2 Circumstellar Disks & Outflows: Observations

3 δ Sco and Be stars as laboratories for CS disk physics

4 Dynamics of Circumstellar Material and tidal interactions in hot binaries

5 Massive star formation out of a dynamic environment

6 Magnetospheres of Hot Stars

VENUE The workshop will take place in Foz do Iguaçu, Brazil, close to the magnificent Iguaçu Waterfalls, a network of 275 waterfalls in the Iguaçu River that lies in the border of Brazil and Argentina. The site was designated World Heritage by UNESCO. Tourist attractions include visits to both the Brazilian and Argentinian sides of the Falls, natural parks, and the dam of the Itaipu Hydroeletric Facility. In 1994, the American Society of Civil Engineers elected the Itaipu Dam as one of the seven modern Wonders of the World.

The Meeting will take place in the Rafain Hotel and Convention Centers (http://www.rafainpalace.com.br/v2/home/). Special rates are available for the period of the conference. Hotel costs are 274 BRL (170 USD) for single occupancy and 171 BRL (106 USD) per person for double occupancy. Those rates include full board (breakfast, morning coffebreak, lunch, afternoon coffebreak and dinner).

IMPORTANT NOTE: all participants are encouraged to register in this hotel for two reasons. First, those low rates will only be secured if a minimum of 50 rooms is booked. Second, the hotel is far from the city (10 km) and there are no restaurants nearby.

PRE-REGISTRATION If you intend to participate in the workshop please fill in the pre-registration form in our web page. This is not a formal registration, and requires no commitment from you. Our goal is to have an idea about the number of participants for organization purposes.

PROCEEDINGS We will have online conference proceedings, whether these will be published in print depends on funding decisions made towards the end of 2011 only.

FINANCIAL MATTERS The workshop fee will be 200 USD. We will have some limited financial support for students and young researchers. Money will be requested for the Brazilian students that do not have research contingency funds.

DEADLINES Requests for financial support must be submitted by Nov. 1^{st} at the latest, together with the abstract of the intended contribution. The deadline for the final registration and abstracts is Dec. 17^{th}.

CONTACT

– Website: http://www.eso.org/sci/meetings/2012/csdyn.html

– Mail: csdyninfo@eso.org

SOC: A.C. Carciofi (co-chair), D. Baade, J.E. Bjorkman, A. Damineli, W. Dent, A. Domiciano de Souza, Th. Rivinius (co-chair), S. Štefl, J. Vink, G. Wade

LOC: A. Carciofi (chair), A. Damineli, M. Borges, M. Teodoro, C. Barbosa, D.M. Faes, M.E. Gomez, Th. Rivinius and C. Martayan

Email: carciofi@usp.br

See also http://www.eso.org/sci/meetings/2012/csdyn.html